# HNN extensions of free groups with equal associated subgroups of finite index: polynomial time word problem

### v1

@@ 1. Introduction
The study of computational problems in the theory of groups began in the early twentieth century. Two central themes in this area are decidability and computational complexity, that together shape our understanding of which problems can be solved algorithmically and how efficiently. In his 1911 work [4], M. Dehn introduced three fundamental decision problems: the word problem, the conjugacy problem, and the isomorphism problem, that have since been central to the field. A significant result concerning decidability was established in the 1960s when Novikov [13] and Boone [1] demonstrated the existence of finitely presented groups for which the word problems are undecidable. Nevertheless, for many important classes of groups, such as automatic groups, finitely generated linear groups, and finitely presented residually free groups, the word problem remains decidable.
@@ 1.1. HNN extensions
Let $G={\left\langle X\mid R\right\rangle}$, $H_{1},H_{2}\leq G$ and $\varphi:H_{1}\rightarrow H_{2}$ be a group isomorphism. The HNN extension of $G$ relative to $\varphi$ is the group denoted by $G\ast_{\varphi}t$, given by the following presentation:
@@
where $w_{0},\dots,w_{k}$ are group-words over the alphabet of $G$, called syllables, and $\varepsilon_{i}=\pm 1$. The number $k$ is called the syllable length of $w$.
@@
  • For ascending HNN extensions (when one of the subgroups is the entire group $G$) Lohrey [10] established polynomial-time decidability using straight-line programs.
@@ 1.2. Our results
The main contributions of this paper are summarized in the following theorems.
@@
Suppose that $H_{1}=H_{2}$ are normal subgroups of $F$ of finite index and let $\varphi:H_{1}\to H_{2}$ be an isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
Theorem 4.3 can be generalized to the case where $H_{1}=H_{2}$ are subgroups of $F$ of finite index and $\varphi$ can be restricted to an isomorphism $\varphi:N\to N$ of a normal subgroup $N\unlhd F$ of finite index. We call such $\varphi$ normalizable in Section 5.
@@
Suppose that $H_{1}=H_{2}$ are subgroups of $F$ of finite index and let $\varphi:H_{1}\to H_{2}$ be a normalizable isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
@@ 1.3. Outline
The paper is organized as follows. Section 2.1 introduces essential preliminaries of free groups and subgroup graphs. In Section 3 we discuss the definition and basic properties of straight-line programs. Section 4 presents a polynomial-time algorithm for the word problem in $F\ast_{\varphi}t$ in the case where $H_{1}=H_{2}$ are normal subgroups of finite index, which establishes Theorem 4.3. Section 5 defines the property of $\varphi:H_{1}\to H_{2}$ to be normalizable, presents a polynomial-time algorithm for the word problem in $F\ast_{\varphi}t$ in the case where $H_{1}=H_{2}$ are subgroups of finite index and $\varphi$ is normalizable, which establishes Theorem 5.15.
@@ 2.2. Subgroup graph
Folded graphs have the following important property: for any path $p$ we have
@@ 2.4. Schreier graph
If $H$ is a subgroup of $F(X)$, then there is a unique (up to an isomorphism) subgroup graph $\Gamma$ satisfying $L(\Gamma)=H$. Denote such graph by $\Gamma_{H}$.
@@ 2.5. Subgroup graph homomorphism
$H_{1}\leq H_{2}\ \Leftrightarrow\$there exists a homomorphism $\varphi:\Gamma_{H_{2}}\to\Gamma_{H_{1}}$.
@@ 2.6. Regularity, self-similarity, and shift operation
Here we introduce the shift operation on subgroup graphs and discuss two properties that allow it to be computed efficiently. We say that a subgroup graph $\Gamma=(V,E^{\pm},\mu,r)$ is $X$-regular (or deterministic) if for each vertex $v$ of $\Gamma$ and for each $x\in X^{\pm}$, there is exactly one edge from $v$ labeled with $x$.
@@
Let $\Gamma$ be the subgroup graph of $H\leq F(X)$. Then
@@
By $\operatorname{{Aut}}(\Gamma)$ we denote the group of automorphisms of $\Gamma$. We say that $\Gamma=(V,E^{\pm},\mu,r)$ is self-similar if for every $u,v\in V$ there is an automorphism of $\Gamma$ that maps $u$ to $v$; such automorphism is unique when exists. Denote that automorphism by $S_{u,v}$. Note that $S_{u,v}$ induces
@@
We use the same notation $S_{u,v}$ for the induced functions. Clearly, shift operators preserve labels, i.e., for every $u,v$ and a sequence of edges $p$ we have
@@ 3.1. Definition of a straight-line program
  • $X=\{x_{1},\ldots,x_{n},\varepsilon\}$ is a finite set of terminal symbols (the alphabet), where $\varepsilon$ is a special symbol that denotes the empty string.
@@
  • $R\in{\mathcal{N}}$ is the root symbol.
  • $\delta:{\mathcal{N}}\to X\cup({\mathcal{N}}\times{\mathcal{N}})$ is a production function that determines the set of production rules. There are two types of production rules defined by $\delta(N)$ for $N\in{\mathcal{N}}$:
    • $\delta(N)=x\in X$,
@@
The production graph for $P$ is a directed graph $G(P)=(V,E)$, where $V=X\sqcup{\mathcal{N}}$ and
$E=$ $\ \ \ \{(N,\delta(N))\mid\delta(N)\in X\}$
@@ 3.2. Data representation for SLPs
  • (Assumption-II). The function $\delta$ is stored in a container that enables constant $O(1)$ time complexity for the following manipulations:
@@
We claim that we can make these assumptions when analyzing polynomial-time complexity. Indeed, let us compare our assumptions to a more realistic implementation for SLPs that defines a non-terminal as a natural number and $\delta$ as a trie that maps natural numbers from $\mathcal{N}$$\subset{\mathbb{Z}}$ (written in binary) to elements in $X\cup({\mathcal{N}}\times{\mathcal{N}})$. For this implementation, the $O(1)$ constant-time bound is replaced by an $O(\log_{2}(|{\mathcal{N}}|))$ bound, and the overall time complexity increases by a factor of $\log_{2}(|{\mathcal{N}}^{\ast}|)$, where $|{\mathcal{N}}^{\ast}|$ denotes the size of the largest SLP used in the computations. Clearly, this choice of defining $\delta$ as a trie preserves the property to be polynomial-time computable.
@@ 3.3. Basic properties
Hence, we can decide if $\operatorname{val}(N)=\varepsilon$ for all non-terminals $N\in{\mathcal{N}}$ in linear time $O(|{\mathcal{N}}|)$ starting from the root $R$. ∎
@@
Clearly, for every $N\in{\mathcal{N}}$ we have
@@
Clearly, we can use these formulae to compute $\operatorname{{first}}(N)$ for all non-terminals $n\in{\mathcal{N}}$ in linear time $O(|{\mathcal{N}}|)$ starting from the root $R$. $\operatorname{{last}}(N)$ can be computed in a similar way. ∎
@@
Clearly, the statement holds when $|w|=0$ or $|w|=1$. Let $X^{\prime}\subseteq X$ be the set of letters involved in $w$. Let ${\mathcal{N}}=X^{\prime}\cup\{A_{1},\dots,A_{k-1}\}$. Define an SLP $P_{w}=(X,{\mathcal{N}},A_{k},\delta)$, where $\delta$ is defined as follows:
@@
  • $\delta(A_{2})=(A_{1},A_{x_{3}}),\dots,\delta(A_{k-1})=(A_{k-1},A_{x_{k}})$.
@@ 3.5. Straight-line program over an $X$-digraph
Clearly, $\operatorname{val}(P)$ is a path if and only if
@@
We say that $P$ is reduced if its label $\operatorname{val}(P)$ is reduced as an element of $F(E)$. To reduce $P$ means to find an SLP $P^{\prime}$ such that $o(P^{\prime})=o(P)$, $\mu(\operatorname{val}(P^{\prime}))=\mu(\operatorname{val}(P))$, and $\operatorname{val}(P^{\prime})$ is reduced.
@@ 4. The case of equal associated normal subgroups of finite index
In this section, we assume that the associated subgroups $H_{1}$ and $H_{2}$ are equal, normal in $F$, and of finite index. Let $\Gamma=(V,E^{\pm},\mu,r)$ be the subgroup graph for $H_{1}=H_{2}$. These assumptions imply that
@@
  • The subgroup graph $\Gamma=(V,E^{\pm},r,\mu)$ for $H_{1}=H_{2}$.
@@ 4.1. Application of $\varphi^{\pm 1}$ to an SLP
Let $P$ be an SLP over $\Gamma$ such that $\operatorname{val}(P)$ is a circuit in $\Gamma$ from $r$ to $r$. Then $\mu(\operatorname{val}(P))\in H$ and $\varphi$ is applicable to $\mu(\operatorname{val}(P))$.
@@
There is an algorithm that for a given $P$ produces an SLP $P^{\prime}$ in $O(|P|)$ time satisfying the following conditions:
@@
(Case-I) For each terminal $e\in T$ the algorithm deletes $e$ and $e^{-1}$ from the definitions of all non-terminals. That effectively deletes all occurrences of $e$ and $e^{-1}$ from $\operatorname{val}(P)$.
(Case-II) For each terminal $e\in E^{+}\setminus T$ perform the following.
@@
  • Add non-terminals $N_{e}$ and $N_{e^{-1}}$ with productions $\delta(N_{e})=P_{e}^{\varphi}$ and $\delta(N_{e^{-1}})=P_{e^{-1}}^{\varphi}$.
  • Delete every non-terminal $N$ with production $\delta(N)=e$ from the description of $P$, and in all other productions replace each occurrence of $N$ with $N_{e}$.
  • Delete every non-terminal $N$ with production $\delta(N)=e^{-1}$ from the description of $P$, and in all other productions replace each occurrence of $N$ with $N_{e^{-1}}$.
This effectively replaces every occurrence of $e$ and $e^{-1}$ in $\operatorname{val}(P)$ with $\operatorname{val}(P_{e}^{\varphi})$ and $\operatorname{val}(P_{e^{-1}}^{\varphi})$, respectively.
By construction, all three properties hold for the obtained $P^{\prime}$. ∎
@@ 4.2. The word problem algorithm
Consider a segment $P_{i-1},t^{-1},P_{i},t,P_{i+1}$ in (5). It requires $O(|P_{i-1}|+|P_{i}|+|P_{i+1}|)$ time to check if $P_{i}$ defines an element in $H_{1}$ (i.e., if $\mu(\operatorname{val}(P_{i}))\in H_{1}$) and, if so, to construct an SLP $P$ satisfying the following properties:
@@
The same holds for segments $P_{i-1},t,P_{i},t^{-1},P_{i+1}$, when $P_{i}$ defines an element in $H_{2}$.
@@
By Lemmas 3.2 and 3.5, we can check if $\operatorname{val}(P_{i})$ defines a path that starts and ends at $r$ (i.e., if $\mu(\operatorname{val}(P_{i}))\in H_{1}$), in linear $O(|P_{i}|)$ time. By Proposition 4.1, in linear $O(|P_{i}|)$ time we can compute an SLP $P_{i}^{\prime}$ satisfying
@@
To create a required SLP $P$, it remains to properly concatenate the paths $\operatorname{val}(P_{i-1})$, $\operatorname{val}(P_{i}^{\prime})$, $\operatorname{val}(P_{i+1})$, which can be done using shift operators. Use Lemma 3.2 and Proposition 3.6 to compute
$v_{1}=t(\operatorname{{last}}(P_{i-1})),\ P_{i}^{\prime\prime}=S_{r,v_{1}}(P_{i}^{\prime})\ \mbox{ and }\ v_{2}=t(\operatorname{{last}}(P_{i}^{\prime\prime})),\ P_{i+1}^{\prime}=S_{r,v_{2}}(P_{i+1}).$
By definition of $P_{i}^{\prime\prime}$ and $P_{i+1}^{\prime}$, concatenation $\operatorname{val}(P_{i-1})\circ\operatorname{val}(P_{i}^{\prime\prime})\circ\operatorname{val}(P_{i+1}^{\prime})$ is a path in $\Gamma$ that starts at $r$. Its label defines the same element as the right-hand side of (b) because shift operators preserve labels. Therefore, the SLP $P=P_{i-1}\circ P_{i}^{\prime\prime}\circ P_{i+1}^{\prime}$ satisfies (a) and (b).
@@
$|P_{i}^{\prime\prime}|\leq|P_{i}^{\prime}|\leq|P_{i}|+\sum_{e\in E^{+}\setminus T}|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|\ \mbox{ and }\ |P_{i+1}^{\prime}|\leq|P_{i+1}|$
because applying shift operators does not increase the size of an SLP. Thus,
@@
Suppose that $H_{1}=H_{2}$ are normal subgroups of $F$ of finite index and let $\varphi:H_{1}\to H_{2}$ be an isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
@@
Consider a word $w$ of type (1). If the syllable length $k$ of $w$ is trivial, then we directly check if $w_{0}$ is trivial in the base group $F$.

### v2

@@ 1. Introduction
The study of computational problems in the theory of groups began in the early twentieth century. Two central themes in this area are decidability and computational complexity, that together shape our understanding of which problems can be solved algorithmically and how efficiently. In his 1911 work [4], M. Dehn introduced three fundamental decision problems: the word problem, the conjugacy problem, and the isomorphism problem, that have since been central to the field. A significant result concerning decidability was established in the 1960s when P. Novikov [13] and W. Boone [1] demonstrated the existence of finitely presented groups for which the word problems are undecidable. Nevertheless, for many important classes of groups, such as automatic groups, finitely generated linear groups, and finitely presented residually free groups, the word problem remains decidable.
@@ 1.1. HNN extensions
Let $G={\left\langle X\mid R\right\rangle}$ be a group, $H_{1},H_{2}\leq G$ and $\varphi:H_{1}\rightarrow H_{2}$ be a group isomorphism. The HNN extension of $G$ relative to $\varphi$ is the group denoted by $G\ast_{\varphi}t$, given by the following presentation:
@@
where $w_{0},\dots,w_{k}$ are group words over the alphabet $X$ of $G$, called syllables, and $\varepsilon_{i}=\pm 1$. The number $k$ is called the syllable length of $w$.
@@
  • For ascending HNN extensions (when one of the subgroups is the entire group $G$) M. Lohrey [10] established polynomial-time decidability using straight-line programs.
@@ 1.2. Our results
Throughout the paper $F$ denotes the free group $F(X)$ over a finite alphabet $X$ and $H$ denotes a finitely generated subgroup of $F$. The main contributions of this paper are summarized in the following theorems.
@@
Suppose that $H$ is a normal subgroup of $F$ of finite index and let $\varphi:H\to H$ be an automorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
Theorem 4.3 can be generalized to the case where $H$ is a subgroup of $F$ of finite index and $\varphi$ can be restricted to an automorphism $\varphi:N\to N$ of a normal subgroup $N\unlhd F$ of finite index. We call such $\varphi$ normalizable in Section 5.
@@
Suppose that $H$ is a subgroup of $F$ of finite index and let $\varphi:H\to H$ be a normalizable isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
@@ 1.3. Outline
The paper is organized as follows. Section 2.1 introduces essential preliminaries of free groups and subgroup graphs. In Section 3 we discuss the definition and basic properties of straight-line programs. Section 4 presents a polynomial-time algorithm for the word problem in the HNN extension $F\ast_{\varphi}t$ with equal associated normal subgroups of finite index, which establishes Theorem 4.3. Section 5 introduces the notion of a normalizable isomorphism $\varphi\colon H\to H$ and presents a polynomial-time algorithm for the word problem in the HNN extension $F\ast_{\varphi}t$ with equal associated subgroups of finite index and normalizable $\varphi$, which establishes Theorem 5.16.
@@ 2.2. Subgroup graph
Folded graphs with inversible edges have the following important property: for any path $p$ we have
@@ 2.4. Schreier graph
If $H$ is a subgroup of $F(X)$, then there is a unique (up to an isomorphism) subgroup graph $\Gamma$ satisfying $L(\Gamma)=H$. Denote this graph by $\Gamma_{H}$.
@@ 2.5. Subgroup graph homomorphism
$H_{1}\leq H_{2}\ \Leftrightarrow\$there exists a subgroup graph homomorphism $\varphi:\Gamma_{H_{2}}\to\Gamma_{H_{1}}$.
@@ 2.6. Regularity, self-similarity, and shift operation
Here we introduce the shift operation on subgroup graphs and discuss two properties that allow it to be computed efficiently. We say that a subgroup graph $\Gamma=(V,E^{\pm},\mu,r)$ is $X$-regular (or deterministic) if for each vertex $v$ of $\Gamma$ and for each $x\in X^{\pm}$, there is exactly one edge $e$ with $o(e)=v$ and $\mu(e)=x$.
@@
Let $\Gamma$ be the subgroup graph of $H\leq F$. Then
@@
By $\operatorname{{Aut}}(\Gamma)$ we denote the group of automorphisms of $\Gamma$. We say that $\Gamma=(V,E^{\pm},\mu,r)$ is self-similar if for every $u,v\in V$ there exists an automorphism of $\Gamma$ mapping $u$ to $v$; for a folded graph $\Gamma$ such an automorphism, when it exists, is unique. We denote this automorphism by $S_{u,v}$ and refer to it as a shift operation. Note that $S_{u,v}$ induces
@@
We use the same notation $S_{u,v}$ for the induced functions. Clearly, shift operations preserve labels, i.e., for every $u,v$ and a sequence of edges $p$ we have
@@ 3.1. Definition of a straight-line program
  • $X=\{x_{1},\ldots,x_{n}\}$ is a finite set of terminal symbols (the alphabet).
@@
  • $R\in{\mathcal{N}}$ is the root symbol, also denoted by $\operatorname{root}(P)$.
  • $\delta:{\mathcal{N}}\to X\cup\{\varepsilon\}\cup({\mathcal{N}}\times{\mathcal{N}})$ is a production function that determines the set of production rules, where $\varepsilon$ is a special symbol that denotes the empty word. There are two types of production rules defined by $\delta(N)$ for $N\in{\mathcal{N}}$:
    • $\delta(N)=x\in X\cup\{\varepsilon\}$,
@@
The production graph for $P$ is a directed graph $G(P)=(V,E)$, where $V=X\sqcup\{\varepsilon\}\sqcup{\mathcal{N}}$ and
$E=$ $\ \ \ \{(N,\delta(N))\mid\delta(N)\in X\cup\{\varepsilon\}\}$
@@
We say that a non-terminal $N\in{\mathcal{N}}$ in $P$ is essential if the following two conditions hold:
  • (a)
    The graph $G(P)$ contains a path from the root $R$ to $N$. Otherwise, the value of $N$ does not contribute to $\operatorname{val}(P)$, and $N$ can be removed from $\mathcal{N}$}.
  • (b)
    Either $\delta(N)\neq\varepsilon$ or $N=R$. Otherwise, $N$ is a non-root intermediate non-terminal with $\operatorname{val}(N)=\varepsilon$ and, hence, can be removed from $\mathcal{N}$} (with an appropriate redefinition of $\delta$).
It is easy to see that all nonessential non-terminals can be removed in $O(|P|)$ time; we refer to this process as the pruning procedure. Throughout the paper we assume that all non-terminals are essential.
@@ 3.2. Data representation for SLPs
  • (Assumption-II). The function $\delta$ is stored in a container that enables $O(1)$ time complexity for the following manipulations:
@@
We emphasize that Assumptions I–II are made solely for convenience. Any standard representation of SLPs (e.g., encoding non-terminals as natural numbers and storing the production function $\delta$ as a dictionary) would introduce at most a polylogarithmic overhead in the running time. Since we are only concerned with polynomial-time computability, these assumptions can be adopted without loss of generality.
@@ 3.3. Basic properties
Hence, we can decide if $\operatorname{val}(N)=\varepsilon$ for all non-terminals $N\in{\mathcal{N}}$ in linear time $O(|{\mathcal{N}}|)$ by starting at the root $R$, proceeding recursively down the tree to the leaves, and then traversing back up to the root. ∎
@@
For every $N\in{\mathcal{N}}$ the following holds:
@@
We can use these formulae to compute $\operatorname{{first}}(N)$ for all non-terminals $n\in{\mathcal{N}}$ in linear time $O(|{\mathcal{N}}|)$ as in Lemma 3.1. The value of $\operatorname{{last}}(N)$ can be computed similarly. ∎
@@
Clearly, the statement holds when $|w|=0$ or $|w|=1$. Let $X^{\prime}\subseteq X$ be the set of letters involved in $w$. Let ${\mathcal{N}}=X^{\prime}\cup\{A_{1},\dots,A_{k-1}\}$. Define an SLP $P_{w}=(X,{\mathcal{N}},A_{k-1},\delta)$, where $\delta$ is defined as follows:
@@
  • $\delta(A_{2})=(A_{1},A_{x_{3}}),\dots,\delta(A_{k-1})=(A_{k-2},A_{x_{k}})$.
@@ 3.5. Straight-line program over an $X$-digraph
For an SLP $P$ with all essential non-terminals, $\operatorname{val}(P)$ is a path if and only if
@@
We say that $P$ is reduced if its label $\operatorname{val}(P)$ is reduced as an element of $F(E)$. To reduce $P$ means to find an SLP $P^{\prime}$ satisfying the following conditions:
  • $\operatorname{val}(P^{\prime})$ is reduced as an element of $F(E)$,
  • $\operatorname{val}(P^{\prime})=_{F(E)}\operatorname{val}(P)$,
  • $o(P^{\prime})=o(P)$ or $\operatorname{val}(P^{\prime})=\varepsilon$.
@@ 4. The case of equal associated normal subgroups of finite index
In this section we consider the HNN extension $F\ast_{\varphi}t$ of the free group $F$ with equal associated subgroups $H$, where $H$ is normal in $F$ and of finite index, and where the extension is defined by an automorphism $\varphi\colon H\to H$. Let $\Gamma=(V,E^{\pm},\mu,r)$ be the subgroup graph for $H$. These assumptions imply that
@@
  • The subgroup graph $\Gamma=(V,E^{\pm},r,\mu)$ for $H$.
@@ 4.1. Application of $\varphi^{\pm 1}$ to an SLP
Let $P$ be an SLP over $\Gamma$ such that $\operatorname{val}(P)$ is a circuit in $\Gamma$ from $r$ to $r$. Then $\mu(\operatorname{val}(P))\in H$ and $\varphi$ is applicable to $\mu(\operatorname{val}(P))$. There is an algorithm that for a given $P$ produces an SLP $P^{\prime}$ in $O(|P|)$ time satisfying the following conditions:
@@
(Case-I) For each terminal $e\in T$ and each non-terminal $N$ such that $\delta(N)=e$ or $\delta(N)=e^{-1}$, the algorithm redefines $\delta(N)$ to be $\varepsilon$. This effectively deletes all occurrences of $e$ and $e^{-1}$ from $\operatorname{val}(P)$.
(Case-II) For each terminal $e\in E^{+}\setminus T$ the algorithm performs the following steps:
@@
  • For each non-terminal $N$ such that $\delta(N)=e$, redefine $\delta(N)$ to be $\operatorname{root}(P_{e}^{\varphi})$
  • For each non-terminal $N$ such that $\delta(N)=e^{-1}$, redefine $\delta(N)$ to be $\operatorname{root}(P_{e^{-1}}^{\varphi})$.
This effectively replaces each occurrence of $e$ and $e^{-1}$ in $\operatorname{val}(P)$ with $\operatorname{val}(P_{e}^{\varphi})$ and $\operatorname{val}(P_{e^{-1}}^{\varphi})$, respectively. Finally, the resulting SLP is pruned to remove all nonessential non-terminals. By construction, the obtained SLP $P^{\prime}$ satisfies all three properties. ∎
@@ 4.2. The word problem algorithm
Consider a segment $P_{i-1},t^{-1},P_{i},t,P_{i+1}$ in (5). It requires $O(|P_{i-1}|+|P_{i}|+|P_{i+1}|)$ time to check if $P_{i}$ defines an element in $H$ (i.e., if $\mu(\operatorname{val}(P_{i}))\in H$) and, if so, to construct an SLP $P$ satisfying the following properties:
@@
The same holds for segments $P_{i-1},t,P_{i},t^{-1},P_{i+1}$, when $P_{i}$ defines an element in $H$.
@@
By Lemmas 3.2 and 3.5, we can check if $\operatorname{val}(P_{i})$ defines a path that starts and ends at $r$ (i.e., if $\mu(\operatorname{val}(P_{i}))\in H$) in $O(|P_{i}|)$ time. By Proposition 4.1, in $O(|P_{i}|)$ time we can compute an SLP $P_{i}^{\prime}$ satisfying
@@
To create a required SLP $P$, it remains to properly concatenate the paths $\operatorname{val}(P_{i-1})$, $\operatorname{val}(P_{i}^{\prime})$, $\operatorname{val}(P_{i+1})$, which can be done using shift operations. Use Lemma 3.2 and Proposition 3.6 to compute
$v=t(\operatorname{{last}}(P_{i-1})),\ P_{i}^{\prime\prime}=S_{r,v}(P_{i}^{\prime})\ \mbox{ and }\ P_{i+1}^{\prime}=S_{r,v}(P_{i+1}).$
Observe that $\operatorname{val}(P_{i}^{\prime\prime})$ is a circuit and $t(\operatorname{{last}}(P_{i}^{\prime\prime}))=v$. Hence, by definition of $P_{i}^{\prime\prime}$ and $P_{i+1}^{\prime}$, the concatenation $\operatorname{val}(P_{i-1})\circ\operatorname{val}(P_{i}^{\prime\prime})\circ\operatorname{val}(P_{i+1}^{\prime})$ is a path in $\Gamma$ that starts at $r$. Its label defines the same element as the right-hand side of (b) because shift operations preserve labels. Therefore, the SLP $P=P_{i-1}\circ P_{i}^{\prime\prime}\circ P_{i+1}^{\prime}$ satisfies (a) and (b).
@@
$|P_{i}^{\prime\prime}|=|P_{i}^{\prime}|\leq|P_{i}|+\sum_{e\in E^{+}\setminus T}|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|\ \mbox{ and }\ |P_{i+1}^{\prime}|=|P_{i+1}|$
because applying shift operations does not change the size of an SLP. Thus,
@@
Suppose that $H$ is a normal subgroup of $F$ of finite index and let $\varphi:H\to H$ be an automorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
@@
Consider a word $w$ of type (1). If the syllable length $k$ of $w$ is zero, then we directly check if $w_{0}$ is trivial in the base group $F$.